OpenStudy (anonymous):

if a ball is thrown with 20m/s and takes 2 minutes what will be displacement?

OpenStudy (anonymous):

This comes under physics, but anyway... You're given velocity (20 m/s) and time (2 minutes). It's VERY IMPORTANT to make sure your velocity and time are in the same units. Here, velocity is in terms of metres PER SECOND. and time is in minute. so make sure you convert those two minutes in terms of seconds. so, velocity = 20 m/s. time = 2 minutes = 2 * 60 sec (there are 60 seconds in one minute) = 120 seconds. We know that velocity=displacement/time. that can be rewritten as displacement = velocity * time. So here displacement = 20 * 120...... you do the math. and dont forget to write the answer in units, or metres in this case.

OpenStudy (anonymous):

and if you just wanted the answer and already knew all that, well too bad. :P

OpenStudy (anonymous):

s = v*t = 20 * 120 = 2400 m/s

OpenStudy (anonymous):

2400 m